# Collaborative and Distributed Bayesian Optimization via Consensus: Showcasing the Power of Collaboration for Optimal Design

### v1

@@ Abstract
Optimal design is a critical yet challenging task within many applications. This challenge arises from the need for extensive trial and error, often done through simulations or running field experiments. Fortunately, sequential optimal design, also referred to as Bayesian optimization when using surrogates with a Bayesian flavor, has played a key role in accelerating the design process through efficient sequential sampling strategies. However, a key opportunity exists nowadays. The increased connectivity of edge devices sets forth a new collaborative paradigm for Bayesian optimization. A paradigm whereby different clients collaboratively borrow strength from each other by effectively distributing their experimentation efforts to improve and fast-track their optimal design process. To this end, we bring the notion of consensus to Bayesian optimization, where clients agree (i.e., reach a consensus) on their next-to-sample designs. Our approach provides a generic and flexible framework that can incorporate different collaboration mechanisms. In lieu of this, we propose transitional collaborative mechanisms where clients initially rely more on each other to maneuver through the early stages with scant data, then, at the late stages, focus on their own objectives to get client-specific solutions. Theoretically, we show the sub-linear growth in regret for our proposed framework. Empirically, through simulated datasets and a real-world collaborative material discovery experiment, we show that our framework can effectively accelerate and improve the optimal design process and benefit all participants.
@@ 1 Introduction
The success of many real-world applications critically depends on trial & error. Often the goal is to manipulate a set of variables, called designs, to achieve a desired outcome or system response. For instance, material scientists perform time-consuming and expensive experiments to determine optimal compositions (Zhang et al.,, 2023) that produce a material with desired properties. Similarly, additive manufacturers must calibrate many design parameters, including laser power, beam diameter, and hatching pattern through trial & error (Buchanan and Gardner,, 2019) so that their product matches its intended shape. Those studies, whether through experimentation or simulations like finite element models, consume resources (e.g., time and budget) that can significantly limit progress.
@@
  • The power of collaboration for material discovery: We perform an experiment for accelerated optimization of sustainable functional hydrogels through collaborative high-throughput material manufacturing workflows. The experiment showcases the ability of collaboration to significantly outperform its non-collaborative counterpart.
The remainder of this paper is organized as follows. In Sec. 2, we present a motivating example that drives this work. We then briefly discuss related literature in Sec. 3. In Sec. 4, we formulate the problem mathematically and provide our proposed collaborative BO framework. Sec. 5 and Sec. 6 showcase the benefits of our approach through several simulation studies and an experiment on accelerated material discovery. We conclude our paper with a brief discussion about open problems in this relatively new field in Sec. 7.
@@ 2 Motivation
Our work is motivated by the optimal material discovery of soft functional materials, specifically hydrogels, which drive new solutions to challenges across food, energy, water, and health applications (Chai et al.,, 2017; Chen et al.,, 2022). In optimal material discovery, scientists often aim to discover the material composition (i.e., the portion of all ingredients) that yields a material with optimal quality. However, progress on optimal discovery for soft functional materials has been relatively limited (Almawash et al.,, 2022). The barrier to progress is due to both the complexity of soft functional materials and the resource-intensive nature of processing and characterizing material performance, quality, function, and processability (Xu et al.,, 2019; Ding et al.,, 2021; Zhang et al.,, 2023). There is a critical need for innovation in the area of accelerated materials discovery, as evidenced by the 2021 federal multi-agency Materials Genome Initiative (MGI) for optimizing, discovering and deploying advanced materials twice as fast and at a fraction of the cost compared to traditional methods. Given the current limitation, our central hypothesis is:
Hypothesis: Collaboration can accelerate the pace of optimal material discovery of hydrogels and yield optimal material function with minimal resource expenditures.
Figure 1: Illustration of collaborative optimal material discovery.
Fig. 1 provides a microcosm of our collaborative solution, which we have tested in Sec. 6. Our experiment features multiple collaborating human agents (chemists) that perform material synthesis and characterization via automated robots based on a specific material composition. Then they will coordinate to decide on their next-to-sample composition. Our goal is to rapidly engineer a high-performance hydrogel by identifying optimal component compositions of the hydrogel through collaboration.
@@ 4.2 Collaborative BO via Consensus
Finally, the consensus constraint provides a naturally distributed approach to solve (3) sequentially. To see this, recall that each client can run $T$ experiments across $T$ iterations and the initial dataset is $\mathcal{D}_{k}^{(0)}$. Then a natural algorithm would take the following form: At iteration $t\in\{0,\ldots,T-1\}$, given $\bm{x}_{\mathcal{C}}^{(t)}$, all clients calculate $\bm{x}_{k}^{(t)\text{new}}=[(\bm{W}^{(t)}\otimes\bm{I}_{D})\bm{x}_{\mathcal{C}}^{(t)}]_{k}$. Now, each client $k$ conducts an experiment at $\bm{x}_{k}^{(t)\text{new}}$, and augments its dataset with the new observation. Each client then fits a surrogate (e.g., $\mathcal{GP}$) using the current dataset $\mathcal{D}_{k}^{(t+1)}$ and accordingly constructs the utility function $U(\hat{y}_{k}(\bm{x}_{k}))$. Then the clients maximize their local utility $\max_{\bm{x}_{k}}\mathbb{E}_{\hat{y}_{k}|\mathcal{D}^{(t+1)}_{k}}\left[U(\hat{y}_{k}(\bm{x}_{k}))\right]$ to obtain a new candidate solution $\bm{x}^{(t)}_{k}$. Finally, all clients send their solutions to each other or a central orchestrator, and the process repeats.
@@ 6 Case Study
Given the demand for the accelerated optimal discovery of sustainable functional materials, the case study was focused on the closed-loop optimization of a composite hydrogel derived from biomaterials. Hydrogels are soft materials composed of polymers and water that support critical technologies across various industries and drive new solutions to challenges across food, energy, water, and health applications (Chai et al.,, 2017). Hydrogels can be formed from sustainable materials, which offer environmental benefits through reducing annual plastic waste and atmospheric $\text{CO}_{2}$ recycling (Zhang et al.,, 2022; Mishbak et al.,, 2019). The hydrogel under consideration is prepared from multi-reducing end alginate (M-Alg), carboxymethyl chitosan sodium salt (CMCS), and acetic acid (AcOH). Our goal was to rapidly engineer a high-performance sustainable hydrogel by identifying the optimal component composition (i.e., amount of each component in the material) of the hydrogel through collaboration. We focused on discovering the composition that exhibits an optimized shear storage modulus by characterizing the absolute phase angle $|\varphi|$ of dynamic-mode milli-cantilever sensors; a relationship established in (Zhang et al.,, 2023).
@@
The experiment involved four collaborating human agents and was done in four stages as shown in Fig. 6. Stage 1-synthesis: the hydrogel was synthesized at initially chosen compositions by the four human agents (Zhai et al.,, 2023). CMCS was synthesized based on the procedure from Zhou et al., (2023). M-Alg was dissolved in deionized water (DIW) to generate a 5 wtstock solution. CMCS was dissolved in DIW to generate a 5 wtsolution. The stock precursor solutions were then diluted to 0.75 wtusing DIW. The diluted precursor solutions were then filtered with a 0.45 µm nylon syringe filter. The composition constraints identified for the composite hydrogel were: 0.16 wt$<$ $\text{c}_{\text{M-Alg}}$ $<$ 0.66 wt%, 0.01 wt$<$ $\text{c}_{\text{CMCS}}$ $<$ 0.45 wt%, and 0.01 wt$<$ $\text{c}_{\text{AcOH}}$ $<$ 10 wt%, and $\text{c}_{\text{M-Alg}}$ + $\text{c}_{\text{CMCS}}$ = 0.66 wt%, where wtis with respect to all material components except for water (i.e., the mass of water in each sample is neglected in the weight percentage calculation for data visualization purposes). The mass of water in each sample is obtained from a consistent total mass basis among all samples. These constraints ensured the samples would form a homogeneous hydrogel and exhibit the same polymer mass. M-Alg and CMCS solutions and AcOH were first combined. Additional DIW was subsequently added to the samples to ensure that the total mass per sample was identical followed by mixing. The hydrogels were then obtained by allowing the crosslinking reaction to proceed for 12 hours in a humid environment. Stage 2-characterization: the output was characterized via robotically-directed sensing using a piezoelectric milli-cantilever (PEMC) (Zhang et al.,, 2023). The PEMC sensor phase angle at resonance was obtained using an impedance analyzer (e5061b; Keysight). The PEMC measurement principle can be found in Cesewski et al., (2020); Singh et al., (2021); Haring et al., (2020). In a typical measurement cycle, the PEMC sensor was first inserted into the hydrogel after reaching a stable baseline and then maintained for 60 seconds. Afterward, the sensor was removed and allowed to establish a new baseline for 60 seconds before characterizing the next sample. Stage 3-consensus: the output from each human agent was shared into a cloud storage location where our consensus approach was performed to find the next composition for each agent to test. Stage 4-next iteration: the next-to-sample compositions were then shared from the cloud to the agents for synthesis and characterization. The process was repeated for 5 iterations. The closed-loop experimentation process was stopped after 5 iterations based on the observed convergence of the sensor response in the experimental group with collaboration, which is consistent with previous studies on closed-loop experimentation.
The experiment involved two experimental groups: a non-collaborative group, where four human agents operated in isolation, and a collaborative group. Each agent had an initial dataset with 5 randomly sampled compositions. These initial datasets were the same across the collaborative and non-collaborative groups to ensure a fair comparison. The results are shown in Fig. 7.
Figure 7: Trend in material shear storage modulus (as indicated by sensor phase angle response $|\varphi|$) vs. iteration without (a) and with (b) collaboration. c) Trend of material storage modulus throughout five sequential rounds of closed-loop experimentation.
We found that the collaborative group was able to significantly outperform the non-collaborative group in terms of the quality of the hydrogel and the speed to achieve desired quality. Figs. 7(a) and 7(b) show the trend of the observed $|\varphi|$ with iterations. It can be seen that a collaborative approach consistently yields higher $|\varphi|$ for all clients compared to a non-collaborative approach. Fig. 7(c) shows that the collaborative group helped each other narrow down the exploration space to regions with large absolute values of phase angle where most of the testing happened (see the concentration of brown dots), whereas the non-collaborative group spent a lot of trial effort in regions associated with compositions of relatively inferior mechanical properties. Indeed, after only 5 iterations (each taking $\approx 13$ hrs), all agents in the collaborative groups were able to identify an optimal composition with an absolute value of phase angle close to ${}^{\circ}$. Whereas the non-collaborative group, on average, was still ${}^{\circ}$ degree below this optimal solution. Note here that an iteration takes a similar time across both groups, as each group will synthesize 4 different hydrogel samples to test at each iteration. This shows that not only can we obtain better results across all participants, but also we obtain these results much faster through collaboration.
@@
This small-scale case study serves as a proof-of-concept showcasing that collaboration in optimal design can reap benefits for the participating entities.
@@ 7 Discussion and Conclusion
In this paper, we propose a collaborative Bayesian optimization framework built upon a consensus mechanism. Our experiments on simulated data and a real-world material discovery case study show that collaboration through our proposed framework can accelerate and improve the optimal design process.
@@ Appendix A Technical Results
By Lemma 2, we have, with probability $1-\delta$ ($\delta\in(0,1)$), ${\texttt{EI}}^{(t)}_{k}(\bm{x})\geq I_{k}^{(t)}(\bm{x})-\sqrt{\beta_{k}^{(t)}}\sigma_{k}^{(t)}(\bm{x})$, where $I_{k}^{(t)}(\bm{x})=(f_{k}(\bm{x})-y^{*(t)}_{k})^{+}$ and $\beta^{(t)}_{k}$ is defined in Lemma 2.. Therefore, we obtain
@@
By Lemma 5, we know $\tau(\cdot)$ is Lipschitz continuous on a bounded domain. Assume the Lipschitz constant is $L_{\tau}$, we obtain
@@
$=|\sigma_{k}^{(t)}(\bm{x}^{(t)}_{k})(\tau(z_{k}^{(t)}(\bm{x}^{(t)}_{k}))-\tau(z_{k}^{(t)}(\bm{x}^{(t)\text{new}}_{k})))+\mathcal{O}(\frac{1}{(\log(2+t))^{0.5+\epsilon}})\tau(z_{k}^{(t)}(\bm{x}^{(t)\text{new}}_{k}))|$
@@
where the last inequality uses the fact that $z=\tau(z)-\tau(-z)$ (Lemma 5).
@@ Appendix B Supporting Lemmas and Some Useful Equations
Therefore, $z_{k}^{(t)}(\bm{x})\geq q_{k}^{(t)}(\bm{x})-\sqrt{\beta_{k}^{(t)}}$. By Lemma 5, we know $\tau(\cdot)$ is a non-decreasing function. Therefore, we have
@@
[…]
The last inequality holds due to $\tau(z)\geq z,\forall z$ (Lemma 5). If $I_{k}^{(t)}(\bm{x})=0,$ then we know $q_{k}^{(t)}(\bm{x})\leq 0$ and $\sigma_{k}^{(t)}(\bm{x})\left(q_{k}^{(t)}(\bm{x})-\sqrt{\beta_{k}^{(t)}}\right)<0$. Therefore, we have $\texttt{EI}^{(t)}_{k}(\bm{x})\geq I_{k}^{(t)}(\bm{x})-\sqrt{\beta_{k}^{(t)}}\sigma_{k}^{(t)}(\bm{x})$. If $I_{k}^{(t)}(\bm{x})>0$, then we also have
@@
using the fact that $\tau(z)\leq 1+z,\forall z\geq 0$ (Lemma 5). When $y^{*(t)}_{k}-\mu_{k}^{(t)}(\bm{x}^{(t)\text{new}}_{k})\leq 0$, we know $\forall z\leq 0,\tau(z)\leq\phi(z)\leq 1$ (Lemma 5). Therefore,
@@
B.1 Properties of $\tau(z)$
We present some useful properties of $\tau(z)$ in the following Lemma.
Lemma 5.
The function $\tau(z)=z\Phi(z)+\phi(z)$ has the following properties:
  • $\tau(z)$ is a non-decreasing function in $\mathbb{R}.$
  • $\tau(z)\geq 0,\forall z\in\mathbb{R}.$
  • $\tau(z)\geq z,\forall z\in\mathbb{R}.$
  • $\tau(z)\leq 1+z,\forall z\geq 0.$
  • $\tau(z)\leq 1,\forall z\leq 0$.
  • $z=\tau(z)-\tau(-z),\forall z\in\mathbb{R}.$
  • $\tau(z)$ is Lipschitz continuous on a bounded domain in $\mathbb{R}$.
Proof.
  • We have $\frac{d\tau(z)}{dz}=\Phi(z)+z\phi(z)-z\phi(z)=\Phi(z)\geq 0,\forall z\in\mathbb{R}$. Therefore, $\tau(z)$ is a non-decreasing function in $\mathbb{R}.$
  • By definition, $\texttt{EI}(\bm{x})=\sigma(\bm{x})\tau(z)\geq 0$. Therefore, $\tau(z)\geq 0,\forall z$.
  • By definition, $\tau(z)=z\Phi(z)+\phi(z)$. If $z\leq 0,$ then it is trivial to show $\tau(z)\geq 0\geq z$. If $z>0$, we want to show $z\Phi(z)+\phi(z)-z\geq 0$. Let $g(z)=z\Phi(z)+\phi(z)-z$. We have $\frac{dg(z)}{dz}=\Phi(z)+z\phi(z)+\frac{d\phi(z)}{dz}-1=\Phi(z)+z\phi(z)-z\phi(z)-1=\Phi(z)-1\leq 0$. Therefore, $\tau(z)-z$ is a non-increasing function. As $z\to\infty$, we have $\lim_{z\to\infty}(\tau(z)-z)=\lim_{z\to\infty}(z\Phi(z)+\phi(z)-z)=z+0-z=0$. Therefore, $\tau(z)-z\geq 0$ and we conclude our proof.
  • When $z\geq 0$, we have $\tau(z)=z\Phi(z)+\phi(z)\leq z+\phi(z)\leq z+1$.
  • When $z\leq 0$, we have $\tau(z)=z\Phi(z)+\phi(z)\leq\phi(z)\leq 1$.
  • $\tau(z)-\tau(-z)=z\Phi(z)+\phi(z)-(-z\Phi(-z)+\phi(-z))=z\Phi(z)+\phi(z)-(-z(1-\Phi(z))+\phi(z))=z$.
  • The CDF $\Phi(z)$ is Lipschitz continuous since it has a bounded derivative. Without loss of generality, assume the Lipschitz constant is $L_{\Phi}$. Now, we need to show $z\Phi(z)$ is Lipschitz. For any $z_{1},z_{2}$ in a bounded interval $[a,b]$, we have
    $|z_{1}\Phi(z_{1})-z_{2}\Phi(z_{2})|$ $\leq|z_{1}\Phi(z_{1})-z_{1}\Phi(z_{2})|+|z_{1}\Phi(z_{2})-z_{2}\Phi(z_{2})|$
    $=|z_{1}(\Phi(z_{1})-\Phi(z_{2}))|+|\Phi(z_{2})(z_{1}-z_{2})|$
    $\leq|z_{1}||\Phi(z_{1})-\Phi(z_{2}|+|\Phi(z_{2})||z_{1}-z_{2}|$
    $\leq L_{\Phi}|z_{1}||z_{1}-z_{2}|+|\Phi(z_{2})||z_{1}-z_{2}|$
    $\leq(L_{\Phi}|z_{1}|+|\Phi(z_{2})|)|z_{1}-z_{2}|.$
    Since $z_{1}$ is bounded in $[a,b]$, we know $|z_{1}|$ is finite. Also, $|\Phi(z_{2})|\leq 1$. As a result, we proved that $z\Phi(z)$ is Lipschitz on a bounded interval. Furthermore, since the derivative of $\phi(z)$ is bounded on $[a,b]$, and given the fact that the summation of two Lipschitz continuous functions is still Lipschitz, we know $\tau(z)$ is Lipschitz on a bounded interval.

### v2

@@ Abstract
Optimal design is a critical yet challenging task within many applications. This challenge arises from the need for extensive trial and error, often done through simulations or running field experiments. Fortunately, sequential optimal design, also referred to as Bayesian optimization when using surrogates with a Bayesian flavor, has played a key role in accelerating the design process through efficient sequential sampling strategies. However, a key opportunity exists nowadays. The increased connectivity of edge devices sets forth a new collaborative paradigm for Bayesian optimization. A paradigm whereby different clients collaboratively borrow strength from each other by effectively distributing their experimentation efforts to improve and fast-track their optimal design process. To this end, we bring the notion of consensus to Bayesian optimization, where clients agree (i.e., reach a consensus) on their next-to-sample designs. Our approach provides a generic and flexible framework that can incorporate different collaboration mechanisms. In lieu of this, we propose transitional collaborative mechanisms where clients initially rely more on each other to maneuver through the early stages with scant data, then, at the late stages, focus on their own objectives to get client-specific solutions. Theoretically, we show the sub-linear growth in regret for our proposed framework. Empirically, through simulated datasets and a real-world collaborative sensor design experiment, we show that our framework can effectively accelerate and improve the optimal design process and benefit all participants.
@@ 1 Introduction
The success of many real-world applications critically depends on trial & error. Often the goal is to manipulate a set of variables, called designs, to achieve a desired outcome or system response. For instance, material scientists perform time-consuming and expensive experiments to determine optimal compositions (Zhang et al., 2023a, ) that produce a material with desired properties. Similarly, additive manufacturers must calibrate many design parameters, including laser power, beam diameter, and hatching pattern through trial & error (Buchanan and Gardner,, 2019) so that their product matches its intended shape. Those studies, whether through experimentation or simulations like finite element models, consume resources (e.g., time and budget) that can significantly limit progress.
@@
  • The power of collaboration for optimal sensor design: We perform an experiment for accelerated optimization of sensor design through collaborative finite element analysis (FEA) workflows. The experiment showcases the ability of collaboration to significantly outperform its non-collaborative counterpart.
The remainder of this paper is organized as follows. In Sec. 2, we present a motivating example that drives this work. We then briefly discuss related literature in Sec. 3. In Sec. 4, we formulate the problem mathematically and provide our proposed collaborative BO framework. Sec. 5 and Sec. 6 showcase the benefits of our approach through several simulation studies and an experiment on accelerated sensor design. We conclude our paper with a brief discussion about open problems in this relatively new field in Sec. 7.
@@ 2 Motivation
Our work is motivated by optimal sensor design. Biosensors are critical bioanalytical technologies that enable the selective detection of target analytes and have broad applications ranging from medical diagnostics, health monitoring, food and water safety, and environmental monitoring (Cesewski and Johnson,, 2020). A biosensor is defined as a device that is based on an integrated biorecognition element and transducer (Thévenot et al.,, 2001). There are two main categories of biosensors: (1) device-based biosensors, and (2) methods-based biosensors (e.g., nanobiosensors) (Johnson and Mutharasan,, 2014). Device-based biosensors are often integrated with microfluidics and exhibit form factors that can be physically integrated with and removed from processes (e.g., thin-film, dip-stick) and preserve the characteristics of the sample to be analyzed. Alternatively, methods-based biosensors based on solutions and suspensions of functionalized particles provide detection by mixing with a sample, and thus, are relatively destructive with respect to the sample. Thus, device-based biosensors are commonly used in a “flow-and-measure” or “dip-and-measure” format. In contrast, methods-based biosensors, such as nanobiosensors, are typically used in a “mix-and-measure” format.
While biosensors have been created for sensitive and selective detection of many target analytes ranging from small molecules, proteins, nucleic acids, cells, biomarkers, and pathogens, there remains a need to further optimize biosensor design (e.g., form factor, functionalization) and performance (e.g., sensitivity, measurement confidence, and speed) to meet the constraints and requirements of industrial and commercial applications (Carpenter et al.,, 2018). Current thrust areas in biosensor design and performance optimization can be categorized as driven by experimentation, simulation, data analytics, or combinations thereof (Selmi et al.,, 2017; Polatoğlu et al.,, 2020; Koveal et al.,, 2022; Zhang et al., 2023b, ; Zhang et al.,, 2024). For example, high-throughput experimentation, sensor arrays, FEA, and data-driven biosensing have been leveraged to improve the understanding, design, and performance of biosensors, particularly device-based biosensors whose design, utility, and performance are often linked to the characteristics of experimental measurement formats, such as microfluidic channel design, flow field parameters (e.g., flow rate), and other parameters of the measurement format (e.g., sample injection time). However, it remains a challenge to optimize biosensor performance, particularly in maximizing the amount of target analyte detected, given the high-dimensionality of the design space associated with biosensor design, functionalization (e.g., concentration of immobilized biorecognition elements), and measurement format parameters. In particular, there remains a need to establish closed-loop self-driving workflows for engineering high-performance biosensors, such as by optimizing the biosensor design, functionalization protocols, and measurement formats that synergize with experimentation, simulation, and machine learning. Given the current limitation, our central hypothesis is:
Hypothesis: Collaboration can accelerate the pace of optimal sensor design and yield optimal design with minimal resource expenditures.
Fig. 1 provides a microcosm of our collaborative solution, which we have tested in Sec. 6. Our experiment features multiple collaborating agents (FEA simulations) that perform biosensor design. Then they will coordinate to decide on their next simulations. Our goal is to rapidly discover the biosensor design and measurement format parameters that find the the maximum amount of captured target analyte.
Figure 1: Illustration of closed-loop biosensor and measurement format design optimization driven by consensus BO-driven simulation.
@@ 4.2 Collaborative BO via Consensus
Finally, the consensus constraint provides a naturally distributed approach to solve (3) sequentially. To see this, recall that each client can run $T$ experiments across $T$ iterations and the initial dataset is $\mathcal{D}_{k}^{(0)}$. Then a natural algorithm would take the following form: At iteration $t\in\{0,\ldots,T-1\}$, given $\bm{x}_{\mathcal{C}}^{(t)}$, all clients calculate $\bm{x}_{k}^{(t)\text{new}}=[(\bm{W}^{(t)}\otimes\bm{I}_{D})\bm{x}_{\mathcal{C}}^{(t)}]_{k}$. Now, each client $k$ conducts an experiment at $\bm{x}_{k}^{(t)\text{new}}$, and augments its dataset with the new observation. Each client then fits a surrogate (e.g., $\mathcal{GP}$) using the current dataset $\mathcal{D}_{k}^{(t+1)}$ and accordingly constructs the utility function $U(\hat{y}_{k}(\bm{x}_{k}))$. Then the clients maximize their local utility $\max_{\bm{x}_{k}}\mathbb{E}_{\hat{y}_{k}|\mathcal{D}^{(t+1)}_{k}}\left[U(\hat{y}_{k}(\bm{x}_{k}))\right]$ to obtain a new candidate solution $\bm{x}^{(t+1)}_{k}$. Finally, all clients send their solutions to each other or a central orchestrator, and the process repeats.
@@ 6 Case Study
A case study on consensus BO-driven closed-loop biosensor design optimization was performed (see Fig. 6). The closed-loop workflow was based on simulation (FEA) guided by a decentralized computing process for collaboration (i.e., consensus BO). Given the demand for device-based biosensors for industrial process analytical technology (e.g., for bioprocess monitoring and control) and health monitoring (e.g., via wearable sensors), the case study was focused on closed-loop optimization of a device-based biosensor. Specifically, our case study focused on optimizing a device-based biosensor with a milli-scale transducer used in a continuous flow format, which is consistent with several types of device-based biosensors, including electrochemical, mechanical, and electromechanical biosensors.
@@
[…]
For laminar flow physics:
  • Surface 2 (inlet): normal inflow velocity = $u_{\text{in}}$.
  • Surface 3 (outlet): static pressure = 0 Pa.
  • Surfaces 1 and 4-6 (top and bottom walls, including the sensing surface): no slip.
Here, no slip represents a traditional no slip boundary condition. The fluid properties were defined by water and obtained from the FEA software’s material library. The initial velocity and pressure fields were zero.
For transport of diluted species physics:
  • Surface 1: $J_{0,c}=-r_{\text{ads}}+r_{\text{des}}$.
  • Surface 2: $c=c_{0}$.
  • Surface 3: $J_{0,c}=\bm{n}\cdot D_{i}\nabla c_{i}$.
  • Suraces 4-6: no flux.
[…]
For surface reaction chemistry:
  • Surface 1: reaction rate $=r_{\text{ads}}-r_{\text{des}}$.
  • Surfaces 2-6: no reaction.
The following surface properties were used: $\rho_{s}=$ density of surface sites and site occupancy number $=1$. The surface diffusion of the bound target analyte was assumed to be zero. The initial concentration of bound target was zero.
The convection-diffusion-reaction model was discretized and solved subject to the aforementioned boundary and initial conditions using a physics-controlled adaptive mesh that contained 373004 (plus 351428 degrees of freedom) in the final mesh for a given combination of inputs $[x_{1},x_{2},x_{3}]=[u_{\text{in}},\rho_{s},\text{GP}_{\text{std}}]$, where $\text{GP}_{\text{std}}$ is the standard deviation of the Gaussian pulse associated with the injected sample. The specific values of $u_{\text{in}},\rho_{s},\text{GP}_{\text{std}}$ examined were systematically selected by traditional BO and the consensus BO model, which are referred to as “non-collaborative” and “collaborative” learning, respectively.
The simulation’s output of interest (i.e., the quantity to be optimized) was the maximum fractional surface coverage of the bound target analyte ($\theta_{\text{max}}$) during the transient binding response, which is a fundamental characteristic of biosensor function that is associated with the performance characteristics of sensitivity, detection limit, dynamic range, and speed.
[…]
The case study involved two groups: a non-collaborative group and a collaborative group, each consisting of three clients. The non-collaborative group used conventional centralized BO algorithms, where each client solves its own problem and there is no communication among clients. In contrast, the collaborative group utilized the CBOC algorithm (Algorithm 1) to collaboratively find optimal solutions. For CBOC, we adopted the leader-driven matrix mentioned in Sec. 4.3. Fig. 7 shows the trend of biosensor performance, specifically maximum amount of captured target species, throughout the iterative closed-loop workflow. As shown in Fig. 7, the self-driving workflow driven by consensus BO was capable of learning an optimized biosensor design and measurement format parameter selection after ten rounds of experimentation as evidenced by the trend of $\theta_{\text{max}}$. Additionally, the average biosensor performance was higher and exhibited lower variance in the collaborative group relative to the non-collaborative group, which highlights the value of leveraging collaboration in closed-loop high-throughput experimentation.
Figure 7: Trend in the biosensor’s maximum fractional surface coverage (i.e., the maximum amount of captured target analyte) vs. iteration number (i.e., sequential rounds of simulation) via BO without and with collaboration after ten iterations.
@@
This case study serves as a proof-of-concept showcasing that collaboration in optimal design can reap benefits for the participating entities.
@@ 7 Discussion and Conclusion
In this paper, we propose a collaborative Bayesian optimization framework built upon a consensus mechanism. Our experiments on simulated data and a real-world sensor design case study show that collaboration through our proposed framework can accelerate and improve the optimal design process.
@@ Appendix A Technical Results
By Lemma 2, we have, with probability $1-\delta$ ($\delta\in(0,1)$), ${\texttt{EI}}^{(t)}_{k}(\bm{x})\geq I_{k}^{(t)}(\bm{x})-\sqrt{\beta_{k}^{(t)}}\sigma_{k}^{(t)}(\bm{x})$, where $I_{k}^{(t)}(\bm{x})=(f_{k}(\bm{x})-y^{*(t)}_{k})^{+}$ and $\beta^{(t)}_{k}$ is defined in Lemma 1. Therefore, we obtain
@@
We know $\tau(\cdot)$ is Lipschitz continuous on a bounded domain. Assume the Lipschitz constant is $L_{\tau}$, we obtain
@@
$\leq|\sigma_{k}^{(t)}(\bm{x}^{(t)}_{k})(\tau(z_{k}^{(t)}(\bm{x}^{(t)}_{k}))-\tau(z_{k}^{(t)}(\bm{x}^{(t)\text{new}}_{k})))+\mathcal{O}(\frac{1}{(\log(2+t))^{0.5+\epsilon}})\tau(z_{k}^{(t)}(\bm{x}^{(t)\text{new}}_{k}))|$
@@
where the last inequality uses the fact that $z=\tau(z)-\tau(-z)$.
@@ Appendix B Supporting Lemmas and Some Useful Equations
Therefore, $z_{k}^{(t)}(\bm{x})\geq q_{k}^{(t)}(\bm{x})-\sqrt{\beta_{k}^{(t)}}$. We know $\tau(\cdot)$ is a non-decreasing function. Therefore, we have
@@
$\geq\sigma_{k}^{(t)}(\bm{x})\left(q_{k}^{(t)}(\bm{x})-\sqrt{\beta_{k}^{(t)}}\right).$
The last inequality holds due to $\tau(z)\geq z,\forall z$. If $I_{k}^{(t)}(\bm{x})=0,$ then we know $q_{k}^{(t)}(\bm{x})\leq 0$ and $\sigma_{k}^{(t)}(\bm{x})\left(q_{k}^{(t)}(\bm{x})-\sqrt{\beta_{k}^{(t)}}\right)<0$. Therefore, we have $\texttt{EI}^{(t)}_{k}(\bm{x})\geq I_{k}^{(t)}(\bm{x})-\sqrt{\beta_{k}^{(t)}}\sigma_{k}^{(t)}(\bm{x})$. If $I_{k}^{(t)}(\bm{x})>0$, then we also have
@@
using the fact that $\tau(z)\leq 1+z,\forall z\geq 0$. When $y^{*(t)}_{k}-\mu_{k}^{(t)}(\bm{x}^{(t)\text{new}}_{k})\leq 0$, we know $\forall z\leq 0,\tau(z)\leq\phi(z)\leq 1$. Therefore,